OpenStudy (fairy733189):

The graphs of f(x) and g(x) are shown below: If f(x) = (x + 7)^2, which of the following is g(x) based on the translation? g(x) = x^2 + 18x + 81 g(x) = x^2 + 10x + 25 g(x) = x^2 + 14x + 47 g(x) = x^2 + 14x + 51

OpenStudy (fairy733189):

@Mehek14

OpenStudy (fairy733189):

let me attach the graph

OpenStudy (fairy733189):

I'm not sure because I don't really understand the question

OpenStudy (kky10997):

a is correct

OpenStudy (fairy733189):

ya but I want to know how to do the problem not just know the answer

OpenStudy (fairy733189):

@mathmale

OpenStudy (mathmale):

If f(x) = (x + 7)^2, which of the following is g(x)? Look carefully at the graph of f(x). Notice that the vertex of this graph is at x=-7 The rule is: given w(x)=(x)^2 (which is represented by a parabola), the graph of z(x)=w(x-b) is the same as the graph of w(x), EXCEPT that it has been translated b units to the right. The graph of q(x)=w(x+c) is the same as that of w(x), except that it's been translated c units to the left. See what you can do with this info.